# UDA: Unsupervised Debiasing Alignment for Pair-wise LLM-as-a-Judge

### v1

@@ Introduction
Large Language Models (LLMs) like GPT-4 and Claude-3 now drive progress across NLP, while their daily use in commerce, education, and entertainment seamlessly integrates them into everyday life (Zhao et al. 2025; Minaee et al. 2025). Rigorous evaluation is therefore indispensable: it guides model optimization and equips users with an empirical basis for selection, ensuring robust deployment in the wild (Chang et al. 2023).
@@
The paradigm of ‘LLM-as-a-judge’ (Zheng et al. 2023) delegates the evaluation to a second large model that acts as an autonomous judge: the judge is prompted to score, rank or critique the outputs of the target model without human labels. Yet this very mechanism introduces systematic bias (Wang et al. 2023; Thakur et al. 2025). In this paradigm, pairwise judgments consistently outperform pointwise scoring in robustness and are widely adopted, for example, Zheng et al. first established this in the LMSys Chatbot Arena (Chiang et al. 2024): GPT-4 achieved markedly higher agreement with human preferences when ranking anonymized responses head-to-head than when assigning absolute scores. Subsequent works (Wang, Zhang, and Choi 2025) corroborate that pairwise ranking reduces intra-model variance; on RewardBench, GPT-4o attains 90.5% accuracy in pairwise mode versus 88.0% in pointwise mode, while smaller models trend in the opposite direction. Different LLM are built and trained in unique ways, causing them to have specific biases, such as a preference for wordy or stylistically similar answers. This results in varied scores and rankings. Prior works show that score disparities exceeding 30% among LLMs on identical benchmarks (Ye et al. 2024).
Most researches (Chang et al. 2023) mainly focus on the improvement of evaluation indicators and methods, but the studies on how to reduce the impact of model preferences during the evaluation process are not in-depth enough. Although recent work calibrates proprietary models on-the-fly via prompt engineering (Zhou et al. 2024), the practice remains ad-hoc and theoretically thin. Ensemble schemes (Wataoka, Takahashi, and Ri 2025) mitigate bias but explode in cost, and Nash-based re-ranking (Liu et al. 2025a) scales poorly. In addition, there is relatively little research on how to use deep learning techniques to dynamically adjust the evaluation process, rather than just directly using the evaluation results provided by large models, in order to reduce the bias from different judge models.
@@
The LLM-as-a-judge approach has been widely adopted for evaluating LLMs due to its low cost and high efficiency (Gu et al. 2025). This method leverages one LLM to assess the outputs of another, providing fine-grained evaluations. For instance, OpenAI’s GPT and Anthropic’s Claude are commonly utilized as judge models to evaluate the performance of other LLMs. However, this approach may introduce significant bias. Different judge models can produce varying results and rankings, as their specific traits, differences in training data, and subjective evaluation criteria can influence the evaluation results.
@@
The Elo rating system (Hunter 2004), traditionally used in chess, has been adapted for evaluating large language models (LLMs). Chatbot Arena (Chiang et al. 2024) introduced an Elo system that assesses chatbots through pairwise comparisons based on user votes. This method offers a dynamic and interactive approach to evaluation, providing a degree of objectivity and comparability. However, it may be influenced by user preferences and the diversity of evaluation questions. Yet, this system still relies on user feedback and faces challenges in scalability and scenario diversity. Gray et al. (2022) combined the ELO system with traditional metric-based evaluations, integrating model rankings with specific metric analyses for a more comprehensive assessment. Despite these advancements, existing ELO evaluation systems in LLMs still struggle with balancing the weight of different evaluation dimensions and criteria.
@@
Preference bias in LLMs is a critical evaluation issue (Wataoka, Takahashi, and Ri 2025). It causes LLMs to overrate certain outputs, especially their own output. Studies (Zheng et al. 2023; Xu et al. 2024) reveal that some LLMs favor their own generations. This bias may step from the model’s training data and architecture. Larger and more capable models often show stronger self-preference. The model’s internal probabilities may favor familiar generation styles, and alignment training artifacts could also contribute. Panickssery, Bowman, and Feng explored the link between self-preference bias and self-recognition ability. As shown in Figure2, it reveals heterogeneous LLM preferences.
@@
Other biases also exist. Some models show position bias, preferring the first or last option in pairwise comparisons. Model performance is influenced by whether real names or anonymous aliases are used (Koo et al. 2024). Large models tend to favor their own responses under anonymity, but this self-preference decreases when real names are shown, while smaller models’ self-preference increases.
@@
Zhou et al. (2024) tackled the bias issue in the “LLM-as-a-judge” approach by quantifying the bias in large models’ predictive distributions and then subtracting it to reduce the bias. However, this method has limitations when evaluating close-source llm. Due to the inaccessibility of predictive logic, it is difficult to calculate the distributional differences between pre-trained and SFT models. Consequently, it resorts to prompt engineering to model surface quality and mitigate bias. Liu et al. (2025a) adopted a game-theoretic perspective, framing evaluation as a three-player game and proposing a Nash equilibrium solution. Yet, the equilibrium ratings lack scalability and cannot be directly applied to many real-world large models. Liu et al. (2025b) tried to optimize ELo rating systems in arena settings, they proposed the am-ELO framework, which uses maximum likelihood estimation (MLE) to adjust the Elo probability function by incorporating annotator abilities, enhancing stability and accuracy from an annotation accuracy perspective.
@@
UDA (Unsupervised Debiasing Alignment) is an annotation-free method that reconciles a set of $M$ heterogeneous LLM judges, $\mathcal{J}$, by dynamically adjusting the Elo update rule for a set of $N$ candidate models, $\mathcal{M}$. Given a prompt $x$, a judge $k\in\mathcal{J}$ expresses its raw binary preference $r_{ij}^{(k)}\in\{0,1\}$ for responses $(a_{i},a_{j})$ from models $i,j\in\mathcal{M}$, which in turn determines the Elo score $R_{i}^{(k)}\in\mathbb{R}$ for each model under that judge. The key idea is to treat the collective consensus ranking of all judges as a pragmatic, unsupervised target without human labels.
@@
A small MLP $f_{\theta}:\mathbb{R}^{6D+9}\rightarrow\mathbb{R}^{3}$ then maps this feature vector to the adaptive parameters:
@@
where $K_{ij}$ is the instance-level K-factor, and $s_{i},s_{j}\in(0,1)$ are soft labels that replace the hard $r_{ij}^{(k)}$ to mitigate preference noise ($s_{i}+s_{j}=1$ after adjustment). All parameters $\theta$ are shared across all judges and prompts.

### v2

@@ Introduction
Large Language Models (LLMs) like GPT-4 and Claude-3 now drive progress across Natural Language Processing (NLP), while their daily use in commerce, education, and entertainment seamlessly integrates them into everyday life (ref1zhao2025surveylargelanguagemodels; ref2minaee2025largelanguagemodelssurvey). Rigorous evaluation is therefore indispensable: it guides model optimization and equips users with an empirical basis for selection, ensuring robust deployment in the wild (ref3chang2023surveyevaluationlargelanguage).
@@
The paradigm of ‘LLM-as-a-judge’ (ref4zheng2023judgingllmasajudgemtbenchchatbot) delegates the evaluation to a second large model that acts as an autonomous judge: the judge is prompted to score, rank or critique the outputs of the target model without human labels. Yet this very mechanism introduces systematic bias (ref5wang2023largelanguagemodelsfair; ref6thakur2025judgingjudgesevaluatingalignment). In this paradigm, pairwise judgments consistently outperform pointwise scoring in robustness and are widely adopted, for example, ref4zheng2023judgingllmasajudgemtbenchchatbot first established this in the LMSys Chatbot Arena (ref20chiang2024chatbotarenaopenplatform): GPT-4 achieved markedly higher agreement with human preferences when ranking anonymized responses head-to-head than when assigning absolute scores. Subsequent works (ref13wang2025improvingllmasajudgeinferencejudgment) corroborate that pairwise ranking reduces intra-model variance; on RewardBench, GPT-4o attains 90.5% accuracy in pairwise mode versus 88.0% in pointwise mode, while smaller models trend in the opposite direction. Different LLM are built and trained in unique ways, causing them to have specific biases, such as a preference for wordy or stylistically similar answers. This results in varied scores and rankings. Prior works show that score disparities exceeding 30% among LLMs on identical benchmarks (ref7ye2024justiceprejudicequantifyingbiases).
Most researches (ref3chang2023surveyevaluationlargelanguage) mainly focus on the improvement of evaluation indicators and methods, but the studies on how to reduce the impact of model preferences during the evaluation process are not in-depth enough. Although recent work calibrates proprietary models on-the-fly via prompt engineering (ref10zhou2024mitigatingbiaslargelanguage), the practice remains ad-hoc and theoretically thin. Ensemble schemes (ref11wataoka2025selfpreferencebiasllmasajudge) mitigate bias but explode in cost, and Nash-based re-ranking (ref12liu2025reevaluatingopenendedevaluationlarge) scales poorly. In addition, there is relatively little research on how to use deep learning techniques to dynamically adjust the evaluation process, rather than just directly using the evaluation results provided by large models, in order to reduce the bias from different judge models.
@@
The LLM-as-a-judge approach has been widely adopted for evaluating LLMs due to its low cost and high efficiency (ref14gu2025surveyllmasajudge). This method leverages one LLM to assess the outputs of another, providing fine-grained evaluations. For instance, OpenAI’s GPT and Anthropic’s Claude are commonly utilized as judge models to evaluate the performance of other LLMs. However, this approach may introduce significant bias. Different judge models can produce varying results and rankings, as their specific traits, differences in training data, and subjective evaluation criteria can influence the evaluation results.
@@
The Elo rating system (ref18ELOpaper), traditionally used in chess, has been adapted for evaluating large language models (LLMs). Chatbot Arena (ref20chiang2024chatbotarenaopenplatform) introduced an Elo system that assesses chatbots through pairwise comparisons based on user votes. This method offers a dynamic and interactive approach to evaluation, providing a degree of objectivity and comparability. However, it may be influenced by user preferences and the diversity of evaluation questions. Yet, this system still relies on user feedback and faces challenges in scalability and scenario diversity. gray2022usingeloratingmetric combined the ELO system with traditional metric-based evaluations, integrating model rankings with specific metric analyses for a more comprehensive assessment. Despite these advancements, existing ELO evaluation systems in LLMs still struggle with balancing the weight of different evaluation dimensions and criteria.
@@
Preference bias in LLMs is a critical evaluation issue (ref11wataoka2025selfpreferencebiasllmasajudge). It causes LLMs to overrate certain outputs, especially their own output. Studies (ref4zheng2023judgingllmasajudgemtbenchchatbot; ref15xu2024prideprejudicellmamplifies) reveal that some LLMs favor their own generations. This bias may step from the model’s training data and architecture. Larger and more capable models often show stronger self-preference. The model’s internal probabilities may favor familiar generation styles, and alignment training artifacts could also contribute. ref16panickssery2024llmevaluatorsrecognizefavor explored the link between self-preference bias and self-recognition ability. As shown in Figure 2, it reveals heterogeneous LLM preferences.
@@
Other biases also exist. Some models show position bias, preferring the first or last option in pairwise comparisons. Model performance is influenced by whether real names or anonymous aliases are used (ref17koo2024benchmarkingcognitivebiaseslarge). Large models tend to favor their own responses under anonymity, but this self-preference decreases when real names are shown, while smaller models’ self-preference increases.
@@
ref10zhou2024mitigatingbiaslargelanguage tackled the bias issue in the “LLM-as-a-judge” approach by quantifying the bias in large models’ predictive distributions and then subtracting it to reduce the bias. However, this method has limitations when evaluating close-source llm. Due to the inaccessibility of predictive logic, it is difficult to calculate the distributional differences between pre-trained and SFT models. Consequently, it resorts to prompt engineering to model surface quality and mitigate bias. ref12liu2025reevaluatingopenendedevaluationlarge adopted a game-theoretic perspective, framing evaluation as a three-player game and proposing a Nash equilibrium solution. Yet, the equilibrium ratings lack scalability and cannot be directly applied to many real-world large models. ref19liu2025amelostableframeworkarenabased tried to optimize ELo rating systems in arena settings, they proposed the am-ELO framework, which uses maximum likelihood estimation (MLE) to adjust the Elo probability function by incorporating annotator abilities, enhancing stability and accuracy from an annotation accuracy perspective.
@@
UDA (Unsupervised Debiasing Alignment) is an annotation-free method that reconciles a set of $M$ heterogeneous LLM judges, $\mathcal{J}$, by dynamically adjusting the Elo update rule for a set of $N$ candidate models, $\mathcal{M}$. Given a prompt $x$, a judge $k\in\mathcal{J}$ expresses its raw binary preference $r_{ij}^{(k)}\in\{0,1\}$ for paired responses $(a_{i},a_{j})$ from models $i,j\in\mathcal{M}$, which in turn determines the Elo score $R_{i}^{(k)}\in\mathbb{R}$ for each model under that judge. The key idea is to treat the collective consensus ranking of all judges as a pragmatic, unsupervised target without human labels.
@@
A small Multilayer Perceptron (MLP) $f_{\theta}:\mathbb{R}^{6D+9}\rightarrow\mathbb{R}^{3}$ then maps this feature vector to the adaptive parameters:
@@
where $K_{ij}$ is the instance-level K-factor, and $s_{i},s_{j}\in(0,1)$ are soft labels that replace the hard $r_{ij}^{(k)}$ to mitigate preference noise ($s_{i}+s_{j}=1$ after softmax normalization). All parameters $\theta$ are shared across all judges and prompts.
@@ Conclusion
Acknowledgments
This work was completed during Yang Zhang’s internship at ZhipuAI. We would like to express sincere gratitude to the ZhipuAI data annotation team, including Zikang Wang, Xiaohan Zhang and others, for their diligent efforts in curating the human-annotated datasets that made this research possible.